Andrea ran and biked for a total of 32 miles in 4 hours. Her average running speed was 5 mph and her average biking speed was 13 mph.
Let x= total hours Andrea ran
Let y= total hours Andrea bikes
Use substitution to solve for x and y.
a) How many hours did Andrea run?
b) how many hours did Andrea bike?

Answer by Boreal(15235)

so total distance on bike was 13y miles
and running is 5x miles
5x+13y=32
x+y=4, so x=4-y
5(4-y)+13y=32
20-5y+13y=32
8y=12
y=1.5 hours on the bike or 19.5 miles
x=2.5 hours running or 12.5 miles